OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6.

OpenStudy (anonymous):

"Choose" from where?

OpenStudy (luigi0210):

Welcome to Openstudy :) Use the point-slope formula: \[y-y_{1}=m(x-x_{1})\] You got the information needed: m=3 x1=2 y1=-4

OpenStudy (anonymous):

y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (luigi0210):

Did you plug everything in and solve?

OpenStudy (anonymous):

what is the answer?!

OpenStudy (luigi0210):

We're not here to give you answers.. but if we were we'd give you the wrong one.

OpenStudy (luigi0210):

This, honestly, is a very simple question to do.

OpenStudy (anonymous):

ok then bye bye